OpenStudy (heart_offire):

Find an equation of the line that has intercepts (1,0) and (0,4). A) y = 4x - 1 B) y = -4x + 4 C) y = -4x + 1 D) y = -4x - 4

OpenStudy (anonymous):

I think its C

OpenStudy (michele_laino):

hint: the slope of the requested line is: \[m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{4 - 0}}{{0 - 1}} = ...?\]